### v1

# Respective Roles of Electron-Phonon and Electron-Electron Interactions in the Transport and Quasiparticle Properties of SrVO

Abstract
The spectral and transport properties of strongly correlated metals, such as SrVO (SVO), are widely attributed to electron-electron (-) interactions, with lattice vibrations (phonons) playing a secondary role. Here, using first-principles electron-phonon (-ph) and dynamical mean field theory calculations, we show that -ph interactions play an essential role in SVO: they govern the electron scattering and resistivity in a wide temperature range down to 30 K, and induce an experimentally observed kink in the spectral function. In contrast, the - interactions control quasiparticle renormalizations and low temperature transport, and enhance the -ph coupling. We clarify the origin of the near temperature dependence of the resistivity by analyzing the - and -ph limited transport regimes. Our work disentangles the electronic and lattice degrees of freedom in a prototypical correlated metal, revealing the dominant role of -ph interactions in SVO.
Introduction.—
Strontium vanadate, SrVO (SVO), is a perovskite oxide widely studied as a prototypical correlated metal [1, 2, 3]. Experiments have measured transport and spectral functions in detail in SVO, owing to advances in growth of clean samples [4, 5] and characterization by angle-resolved photoemission spectroscopy [6, 7, 8].
There are clear spectroscopic signatures of strong electron interactions in SVO, including kinks in the quasiparticle dispersion [9, 8] and mass enhancement with quasiparticle weight [6].
In addition, transport measurements have found a near -dependent resistivity in broad temperature ranges below 300 K [10, 11, 12, 13].
These findings are often attributed to strong electron-electron (-) interactions.
As a result, SVO serves as a testbed for theoretical methods treating strongly correlated materials, including first-principles variants of dynamical mean field theory (DMFT) such as density functional theory (DFT)+DMFT [14], GW+DMFT [15, 16, 17, 18], and linear response DMFT [19], and the dynamical cluster [20] and dynamical vertex approximations [21].
However, one can question whether the transport properties and spectral features observed in SVO are the result of purely electronic interactions.
In particular, electron-phonon (-ph) interactions may also play a role in SVO, as they do in other correlated metals where experiments [22, 23, 24] and theory [25, 26, 27] have highlighted the importance of -ph coupling for spectral kinks [22, 28] and electronic transport [29].
A quantitative study combining - and -ph interactions in SVO is needed to clarify the microscopic origin of its electronic behavior.
In this Letter, we show calculations of spectral and transport properties in SVO combining first-principles -ph interactions with DFT+DMFT - interactions [29].
We find that -ph interactions govern the resistivity and its temperature dependence above 30 K,
and account for the experimentally observed kinks
and for most of the linewidth broadening of the spectral functions.
In contrast, the - interactions control the resistivity below 20 K, and are responsible for the most of the quasiparticle mass renormalization.
We also find that the - interactions lead to an enhancement of the effective
-ph coupling.
Our results provide a blueprint for quantifying electronic and lattice contributions to the properties of correlated metals.
Electronic Structure and Electron-Phonon Coupling.—
We calculate the electronic structure, phonon dispersions, and -ph coupling using DFT and density functional perturbation theory (DFPT) with the Quantum Espresso package [30, 31, 32, 33]. We use the experimental lattice parameter of 3.842 Å [2, 13] and project the electronic structure onto the t -orbitals of vanadium [34].
We use Perturbo to compute the -ph interactions, -ph self-energy, spectral functions, and transport [35].
The - self-energy is obtained with DFT+DMFT using the TRIQS code with a continuous-time quantum Monte Carlo solver [36, 37, 38, 39, 40] and Padé analytical continuation [37].
Additional computational details are provided in the Supplemental Material (SM) [41].
As shown in Fig. 1(a), our DFT calculations predict a bandwidth of 2.5 eV for the t electronic bands, which is renormalized by a factor by DMFT, in agreement with experiments [6] and previous DMFT results [14, 16]. In the temperature range we study (115390 K), the imaginary part of the - self-energy, , follows a Fermi liquid behavior.
Figure 1(b) shows that within 100 meV of the Fermi energy can be fit closely by a Fermi liquid parameterization [42],
with 111In the following, we set .
Therefore, based on the Kramers-Kronig relations, and the quasiparticle dispersion near the Fermi energy depend weakly on temperature.
Figure 1(c) shows the phonon dispersions in SVO computed with DFPT and color-coded according to the -ph coupling strength , for each phonon mode and momentum , averaged over the Fermi surface (see SM [41]).
The -ph coupling is stronger for the six highest-energy modes,
which involve distortions of the VO octahedra, such as Jahn-Teller modes.
Spectral Properties.—
We investigate the contributions of -ph and - interactions in SVO by computing the corresponding self-energies [29].
The real and imaginary parts of the - and -ph self-energies at 115 K are shown in Fig. 2(a)-(b), respectively.
The - interactions dominate quasiparticle renormalization, as seen from the greater derivative of compared to within 150 meV of the Fermi energy.
Accordingly, extracting quasiparticle weights with a fit near the Fermi surface, gives a weak contribution to renomalization for -ph interactions () and a dominant contribution for - interactions, with and .
The imaginary part of the self-energy shows an opposite behavior: is much greater than , and thus the -ph interactions account for the majority of electron scattering and spectral width at low energy.
The dominant role of - interactions on quasiparticle renormalization in SVO, despite their small effect on low energy scattering, can be rationalized using the Kramers-Kronig relations [42, 44]:
due to the larger energy scales involved, the - interactions dominate the imaginary part of the self-energy at higher energies (see SM [41]), leading to a greater magnitude (and energy derivative) of compared to at low energy.
We compute the spectral function from the Green’s function
at energy for electron band and momentum . Here, is the DFT band energy, is the Fermi energy, and is the electron self-energy. Following our previous work [29], in separate calculations we compute this Green’s function using the self-energy from DMFT - interactions, the lowest-order self-energy from -ph interactions, and their sum 222The -ph self-energy calculation uses the DFT electron propagator. Unlike in SrRuO [29], using the DMFT electron propagator has a negligible effect in SVO., obtaining corresponding spectral functions capturing different combinations of interactions [Fig. 2(c)-(e)].
The spectral functions from -ph, and those from -ph plus - interactions, show a kink around 60 meV from the Fermi energy which has been observed in experiments [9, 8]. There is a corresponding sharp change in the derivative of at this energy [Fig. 2(a)], whereas this feature is absent in . This result shows that the 60 meV kink observed experimentally in SVO is caused by -ph interactions.
Transport.—
Numerous experiments have measured a near temperature dependence of the resistivity in SVO below 300 K [46, 47, 2, 4, 11, 48, 49, 50, 13].
Due to the strong electronic correlations in SVO, several studies have attributed this resistivity to - interactions in the Fermi liquid regime [46, 11, 50], where behavior is expected. An exception is recent work by Mirjolet et al., who argued that the temperature dependence is better explained by -ph limited resistivity with strong coupling to a dominant phonon mode [12].
Recently, the growth of ultra-clean samples has enabled detailed measurements of the resistivity with reduced defect scattering [4, 13]. In these samples, Ahn et al. [13] find a near- resistivity below 25 K and between about 100300 K, together with a stronger than temperature dependence at intermediate temperatures.
To understand the microscopic origin of this behavior, we compute the resistivity arising from -ph and - interactions using the Green-Kubo formula [51, 29]:
where is the resistivity tensor, and are Cartesian directions, is the energy derivative of the Fermi occupation factor, is the band velocity, and is the spectral function.
The resistivity for different combinations of interactions is shown in Fig. 3(a) and compared with experimental data [11, 13].
Surprisingly, we find that the resistivity is governed by the -ph interactions in SVO. The -ph limited resistivity is an order of magnitude greater than the - limited resistivity, with the latter accounting for only 10% of the experimental value. This result is in contrast with the conventional wisdom that transport properties in SVO are governed by purely electronic interactions. In addition, the contributions are opposite to another prototypical strongly correlated metal, , where the -ph interactions account for only 10% of the resistivity [29].
In SVO, the -ph contribution is similar in magnitude to the experimental value, and the total resistivity including both interactions is in good agreement with experiments.
Below, we show that taking into account the electron correlation induced enhancement of the
-ph interactions increases the resistivity and brings the results in even better agreement with experiments.
The temperature dependence of the -ph limited resistivity is analyzed in more detail in Fig. 3(b), where we show our results on a log-log plot and compare them with experiments [11, 13]. In that plot, the resistivity is computed with the full (iterative) solution of the Boltzmann transport equation (BTE) [35] to include backscattering and improve the treatment of acoustic phonons. The computed -ph limited resistivity follows a temperature dependence between 100200 K, in excellent agreement with the dependence found in experiments in that temperature range [11, 13], and falls to at 300 K. We identify the origin of this nearly temperature trend of the -ph limited resistivity by analyzing the contribution of different phonon modes. Our calculations show that the increasing contribution of strongly coupled optical phonons at higher temperatures is responsible for the dependence of the resistivity between 100200 K (see SM [41]).
Next, we focus on transport at low temperature, where the -ph contribution is expected to be weaker. While DMFT calculations become difficult at low temperatures, we obtain the - limited resistivity by extrapolating our higher-temperature DMFT calculations with a fit. This approach is justified because the - scattering is in the Fermi liquid regime below at least 400 K.
Figure 3(c) shows the computed -ph and DMFT - limited resistivities below 100 K.
We find a clear crossover between 2030 K from -ph to - dominated transport. The -ph limited resistivity becomes much smaller than the - limited resistivity below 20 K, showing that - scattering governs transport at low temperature. This result indicates that - interactions are the origin of the resistivity observed experimentally below 25 K, with the -ph contribution causing deviations from a behavior above 25 K. Note that our DMFT resistivity underestimates the experimental value below 25 K by a factor of 3. We attribute this discrepancy to limitations in the description of electronic interactions in the single site DMFT impurity model [52, 29]. A more accurate treatment may require going beyond single-site DMFT [53, 54].
Correlation-corrected electron-phonon interactions.—
Strong electronic interactions are known to enhance -ph coupling in correlated metals.
For example, calculations using hybrid functionals and the method found correlation-enhanced -ph coupling in unconventional superconductors, attributing the enhancement to decreased electronic screening [25]. Similarly, in multiband -electron systems such as FeSe, treating correlations with DMFT enhances -ph coupling, in this case by increasing the orbital polarization response to phonon perturbations [26].
To study the role of correlations in SVO, we compute the -ph interactions using Hubbard-corrected DFPT (DFPT+) [33, 32], which captures the strong local interactions between -orbitals in a static approximation and accounts for the resulting change in orbital polarization response. We use a Hubbard- parameter of 3 eV, which provides orbital polarization and band structure responses to phonon perturbations similar to a calculation including accurate values of both Hubbard- and Hund’s interactions (see SM [41]).
Adding the Hubbard correction has a small effect on the phonon dispersions, but it enhances the -ph interactions, as shown in Fig. 4(a). The enhancement is mode-dependent and is generally higher for strongly-coupled optical phonons involving VO distortions.
The enhancement is also higher for phonons with momenta away from the point, suggesting a more important role of correlations for distortions breaking lattice-translation symmetry. The spectral and transport properties computed with enhanced -ph coupling from DFPT+ give results qualitatively similar to those from DFPT [41], but with stronger -ph effects.
Notably, the -ph limited resistivity increases by 35% at room temperature [Fig. 4(b)], bringing the resistivity computed with both -ph and - interactions into very good agreement with experiments. For example, the computed resistivity at 290 K is 16 cm, versus an experimental value of 1925 cm [11, 13].
Discussion.—
The origin of the temperature dependence of the resistivity merits further discussion. While the trend for - interactions is expected based on Fermi liquid theory [51], the origin of the near- behavior of the -ph limited resistivity is less clear.
At very low temperatures, the -ph limited resistivity in metals is expected to exhibit a temperature dependence [55] when scattering is dominated by acoustic phonons with momentum .
In our calculations, we find a -ph limited resistivity below 30 K, but the overall resistivity becomes - limited in this temperature range, explaining the experimental resistivity below 25 K.
In the high-temperature limit, based on the temperature dependence of the phonon occupations, one expects a -linear -ph limited resistivity [55].
However, this requires that all phonon modes contribute equally to -ph scattering, with no mode frozen out. While our computed -ph limited resistivity becomes nearly -linear well above 300 K, it is close to a behavior between 100200 K, in agreement with experiments.
As discussed above, this trend is due to the increasing contribution of higher-energy optical phonons with strong -ph coupling for increasing temperatures [41].
Note also that -ph scattering above 50 K in SVO involves phonons with all momenta, ruling out momentum-dependent mechanisms resulting in behavior [56].
Finally, we analyze two approximations made in the -ph transport calculations (see results in SM [41]). First, we examine the use of the lowest-order -ph self-energy [57, 58] by computing the resistivity with a cumulant diagram-resummation method capable of treating delocalized polarons [59].
Including polaron effects leads to a small increase in the resistivity, showing that lowest-order -ph interactions are adequate to describe SVO. Second, we examine the effect of vertex corrections to the current-current correlation function in the Green-Kubo formalism [51, 42].
Vertex corrections improve the description of backscattering and the momentum dependence of -ph scattering, which is particularly important at low temperature [51].
We assess their role above 100 K in the semiclassical limit by comparing the full solution of the BTE, which includes vertex corrections, to the relaxation time approximation, which neglects them.
We find that vertex corrections in the BTE give only a small increase in the resistivity and its temperature dependence. This analysis shows that higher-order -ph interactions and vertex corrections play a minor role in SVO and do not affect our conclusions.
Conclusion.—
In summary, we have shown that -ph interactions play an essential role in
the transport and spectral properties of a prototypical correlated metal, SVO.
In this material, electronic correlations control
other aspects of the low-energy physics,
including the quasiparticle mass renormalization and transport at low temperature. We also found that electronic correlations lead to an effective enhancement of the -ph interactions.
This suggests that SVO may serve as a testbed for investigating the interplay between electron correlations and -ph interactions.
Our results highlight the potential of first-principles calculations combining - and -ph interactions in a consistent way as an emerging tool to study correlated materials.
This work paves the way for a quantitative description of transport and spectral properties
in broad classes of correlated quantum materials.
Acknowledgements.— We thank Andrew Millis and Jennifer Coulter for helpful discussions. D.J.A. is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. This work was also supported by the National Science Foundation under Grant No. OAC-2209262, which provided for code development. D.J.A. and M.B. were partially supported by the AFOSR and Clarkson Aerospace under Grant No. FA95502110460. J.-J.Z. acknowledges support from the National Key R&D Program of China (Grant No. 2022YFA1403400). J.M. is supported by the Slovenian Research Agency (ARIS) under Grants No. P1-0044 and J1-2458. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0026831. The Flatiron Institute is a division of the Simons Foundation.

### v2

@@
-Respective Roles of Electron-Phonon and Electron-Electron Interactions in the Transport and Quasiparticle Properties of SrVO
+Respective Roles of Electron-Phonon and Electron-Electron Interactions in the Transport and Quasiparticle Properties of SrVO3
 Abstract
-The spectral and transport properties of strongly correlated metals, such as SrVO (SVO), are widely attributed to electron-electron (-) interactions, with lattice vibrations (phonons) playing a secondary role.
+The spectral and transport properties of strongly correlated metals, such as SrVO3 (SVO), are widely attributed to electron-electron (-) interactions, with lattice vibrations (phonons) playing a secondary role.
 Here, using first-principles electron-phonon (-ph) and dynamical mean field theory calculations, we show that -ph interactions play an essential role in SVO: they govern the electron scattering and resistivity in a wide temperature range down to 30 K, and induce an experimentally observed kink in the spectral function.
@@
 Introduction.—
-Strontium vanadate, SrVO (SVO), is a perovskite oxide widely studied as a prototypical correlated metal [1, 2, 3].
+Strontium vanadate, SrVO3 (SVO), is a perovskite oxide widely studied as a prototypical correlated metal [1, 2, 3].
 Experiments have measured transport and spectral functions in detail in SVO, owing to advances in growth of clean samples [4, 5] and characterization by angle-resolved photoemission spectroscopy [6, 7, 8].
 There are clear spectroscopic signatures of strong electron interactions in SVO, including kinks in the quasiparticle dispersion [9, 8] and mass enhancement with quasiparticle weight [6].
-In addition, transport measurements have found a near -dependent resistivity in broad temperature ranges below 300 K [10, 11, 12, 13].
+In addition, transport measurements have found a near -dependent resistivity in broad temperature ranges below 300 K [10, 11, 12, 13, 14].
 These findings are often attributed to strong electron-electron (-) interactions.
@@
 We calculate the electronic structure, phonon dispersions, and -ph coupling using DFT and density functional perturbation theory (DFPT) with the Quantum Espresso package [31, 32, 33, 34].
-We use the experimental lattice parameter of 3.842 Å [2, 13] and project the electronic structure onto the t -orbitals of vanadium [34].
+We use the experimental lattice parameter of 3.842 Å [2, 13] and project the electronic structure onto the t2g -orbitals of vanadium [35].
 We use Perturbo to compute the -ph interactions, -ph self-energy, spectral functions, and transport [36].
-The - self-energy is obtained with DFT+DMFT using the TRIQS code with a continuous-time quantum Monte Carlo solver [36, 37, 38, 39, 40] and Padé analytical continuation [37].
+The - self-energy is obtained with DFT+DMFT using the TRIQS code with a continuous-time quantum Monte Carlo solver [37, 38, 39, 40, 41, 42] and Padé analytical continuation [38].
+We use Hubbard-Kanamori interactions with eV and eV to produce spectra in agreement with experiment [6, 7, 8].
 Additional computational details are provided in the Supplemental Material (SM) [43].
 As shown in Fig.
-1(a), our DFT calculations predict a bandwidth of 2.5 eV for the t electronic bands, which is renormalized by a factor by DMFT, in agreement with experiments [6] and previous DMFT results [14, 16].
+1(a), our DFT calculations predict a bandwidth of 2.5 eV for the t2g electronic bands, which is renormalized by a factor by DMFT, in agreement with experiments [6] and previous DMFT results [15, 17].
 In the temperature range we study (115390 K), the imaginary part of the - self-energy, , follows a Fermi liquid behavior.
@@
 The -ph coupling is stronger for the six highest-energy modes,
-which involve distortions of the VO octahedra, such as Jahn-Teller modes.
+which involve distortions of the VO6 octahedra, such as Jahn-Teller modes.
 Spectral Properties.—
@@
 The imaginary part of the self-energy shows an opposite behavior: is much greater than , and thus the -ph interactions account for the majority of electron scattering and spectral width at low energy.
-The dominant role of - interactions on quasiparticle renormalization in SVO, despite their small effect on low energy scattering, can be rationalized using the Kramers-Kronig relations [42, 44]:
+The dominant role of - interactions on quasiparticle renormalization in SVO, despite their small effect on low energy scattering consistent with a Fermi liquid, can be rationalized using the Kramers-Kronig relations [44, 46]:
 due to the larger energy scales involved, the - interactions dominate the imaginary part of the self-energy at higher energies (see SM [43]), leading to a greater magnitude (and energy derivative) of compared to at low energy.
@@
 Following our previous work [30], in separate calculations we compute this Green’s function using the self-energy from DMFT - interactions, the lowest-order self-energy from -ph interactions, and their sum 222The -ph self-energy calculation uses the DFT electron propagator.
-Unlike in SrRuO [29], using the DMFT electron propagator has a negligible effect in SVO., obtaining corresponding spectral functions capturing different combinations of interactions [Fig.
+Unlike in Sr2RuO4 [30], using the DMFT electron propagator has a negligible effect in SVO., obtaining corresponding spectral functions capturing different combinations of interactions [Fig.
 2(c)-(e)].
@@
 Transport.—
-Numerous experiments have measured a near temperature dependence of the resistivity in SVO below 300 K [46, 47, 2, 4, 11, 48, 49, 50, 13].
+Numerous experiments have measured a near temperature dependence of the resistivity in SVO below 300 K [48, 49, 2, 4, 11, 50, 51, 52, 13, 14].
 Due to the strong electronic correlations in SVO, several studies have attributed this resistivity to - interactions in the Fermi liquid regime [48, 11, 52], where behavior is expected.
 An exception is recent work by Mirjolet et al., who argued that the temperature dependence is better explained by -ph limited resistivity with strong coupling to a dominant phonon mode [12].
-Recently, the growth of ultra-clean samples has enabled detailed measurements of the resistivity with reduced defect scattering [4, 13].
+Recently, the growth of ultra-clean samples has enabled detailed measurements of the resistivity with reduced defect scattering [4, 13, 14].
 In these samples, Ahn et al.
+[13] and Brahlek et al.
 [14] find a near- resistivity below 25 K and between about 100300 K, together with a stronger than temperature dependence at intermediate temperatures.
@@
 where is the resistivity tensor, and are Cartesian directions, is the energy derivative of the Fermi occupation factor, is the band velocity, and is the spectral function.
-The resistivity for different combinations of interactions is shown in Fig. 3(a) and compared with experimental data [11, 13].
+The resistivity for different combinations of interactions is shown in Fig. 3(a) and compared with experimental data [11, 13, 14].
 Surprisingly, we find that the resistivity is governed by the -ph interactions in SVO.
@@
 -ph interactions increases the resistivity and brings the results in even better agreement with experiments.
-The temperature dependence of the -ph limited resistivity is analyzed in more detail in Fig.
-3(b), where we show our results on a log-log plot and compare them with experiments [11, 13].
+The temperature dependence of the -ph limited resistivity is analyzed in more detail in Fig. 3(b), where we show our results on a log-log plot and compare them with experiments [11, 13, 14].
 In that plot, the resistivity is computed with the full (iterative) solution of the Boltzmann transport equation (BTE) [36] to include backscattering and improve the treatment of acoustic phonons.
-The computed -ph limited resistivity follows a temperature dependence between 100200 K, in excellent agreement with the dependence found in experiments in that temperature range [11, 13], and falls to at 300 K.
+The computed -ph limited resistivity follows a temperature dependence between 100200 K, in excellent agreement with the dependence found in experiments in that temperature range [11, 13, 14], and falls to at 300 K.
 We identify the origin of this nearly temperature trend of the -ph limited resistivity by analyzing the contribution of different phonon modes.
@@
 This result indicates that - interactions are the origin of the resistivity observed experimentally below 25 K, with the -ph contribution causing deviations from a behavior above 25 K.
-Note that our DMFT resistivity underestimates the experimental value below 25 K by a factor of 3.
-We attribute this discrepancy to limitations in the description of electronic interactions in the single site DMFT impurity model [52, 29].
-A more accurate treatment may require going beyond single-site DMFT [53, 54].
+Note that our DMFT resistivity underestimates the experimental value below 25 K by a factor of 2-3.
+We attribute this discrepancy to limitations in the single site DMFT impurity model, as transport can be sensitive to the methods used to downfold orbitals and effective interactions [54, 55, 56, 30], the lack of non-local contributions to the DMFT self-energy [57], and corrections to the current-current vertex [58, 59, 60, 61].
 Correlation-corrected electron-phonon interactions.—
-Strong electronic interactions are known to enhance -ph coupling in correlated metals.
+Strong electronic interactions are known to significantly modify -ph interactions [62, 63, 64, 65]. In correlated metals, -ph coupling is often enhanced.
 For example, calculations using hybrid functionals and the method found correlation-enhanced -ph coupling in unconventional superconductors, attributing the enhancement to decreased electronic screening [26].
 Similarly, in multiband -electron systems such as FeSe, treating correlations with DMFT enhances -ph coupling, in this case by increasing the orbital polarization response to phonon perturbations [27].
-To study the role of correlations in SVO, we compute the -ph interactions using Hubbard-corrected DFPT (DFPT+) [33, 32], which captures the strong local interactions between -orbitals in a static approximation and accounts for the resulting change in orbital polarization response.
-We use a Hubbard- parameter of 3 eV, which provides orbital polarization and band structure responses to phonon perturbations similar to a calculation including accurate values of both Hubbard- and Hund’s interactions (see SM [41]).
+To study the role of correlations in SVO, we compute the -ph interactions using Hubbard-corrected DFPT (DFPT+) [34, 33].
+DFT+ can be viewed as a static approximation of DFT+DMFT [66]; DFT+DMFT -ph coupling calculations are not currently possible.
+Like DFT+DMFT, DFT+ captures the strong local interactions between -orbitals and the resulting change in orbital polarization response, which modifies -ph coupling.
+We use a Hubbard- parameter of 3 eV in the atomic orbital basis, which is somewhat empirical but provides orbital polarization and band structure responses to phonon perturbations similar to a calculation including a Hubbard- parameter from linear response and a reasonable value of Hund’s .
+The orbital occupation responses are also similar to those found using DMFT with the same orbitals and interaction parameters used throughout (see SM [43])
 Adding the Hubbard correction has a small effect on the phonon dispersions, but it enhances the -ph interactions, as shown in Fig.
 4(a).
-The enhancement is mode-dependent and is generally higher for strongly-coupled optical phonons involving VO distortions.
+The enhancement is mode-dependent and is generally higher for strongly-coupled optical phonons involving VO6 distortions.
 The enhancement is also higher for phonons with momenta away from the point, suggesting a more important role of correlations for distortions breaking lattice-translation symmetry.
@@
 4(b)], bringing the resistivity computed with both -ph and - interactions into very good agreement with experiments.
-For example, the computed resistivity at 290 K is 16 cm, versus an experimental value of 1925 cm [11, 13].
+For example, the computed resistivity at 290 K is 16 cm, versus an experimental value of 1925 cm [11, 13, 14].
 Discussion.—
@@
 in broad classes of correlated quantum materials.
-Acknowledgements.— We thank Andrew Millis and Jennifer Coulter for helpful discussions.
+Acknowledgements.— We thank Andrew Millis, Jennifer Coulter, and Roman Engel-Herbert for helpful discussions.
 D.J.A.